OpenStudy (anonymous):

Given log 3=0.4471, log 5=0.699, log7=0.8451, log2=0.3010, find the value of 1.log 2.4

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Oh logarithms, how they delight me :D

OpenStudy (anonymous):

:W

terenzreignz (terenzreignz):

Can you express 2.4 as a product/quotient of 3's, 2's and 5's ?

OpenStudy (anonymous):

uhm,, i really don't understand,, :O

terenzreignz (terenzreignz):

This takes creativity and/or getting used to. \[\Large 2.4 = \frac{3\cdot 2^2}{5}\]

OpenStudy (anonymous):

2.4 = 24/10, so log (2.4) = log (24/10) = log 24 - log 10 (we know that log 10 = 1) All what we need is the log 24. log (24) = log (2 x 2 x 2 x 3) = log 2+ log 2 + log 2 + log 3 = 0.3010 + 0.3010 + 0.3010 + 0.4471 = 1.3501

OpenStudy (anonymous):

@terenzreignz ,, how'd you come up with that?

OpenStudy (anonymous):

You could have also said that log (24) = log(2^3 x 3) = log (2^3) + log 3 = 3log 2 + log 3 = 1.3501

terenzreignz (terenzreignz):

experience... it's expressing 2.4 in terms of numbers that are useful to you, in this case, only 2's 3's 5's and 7's

OpenStudy (anonymous):

okaaay,, then what??

OpenStudy (anonymous):

See above.

terenzreignz (terenzreignz):

Well, get the log. \[\Large \log(2.4) = \log\left(\frac{3\cdot 2^2}{5}\right)\]

terenzreignz (terenzreignz):

And then use the properties of logs, namely: \[\Large \log(MN) = \log(M) + \log(N)\]\[\Large \log\left(\frac{M}{N}\right)=\log(M) - \log(N) \qquad N\ne 0\]\[\Large \log(M^p) = p\log(M)\]

OpenStudy (anonymous):

wait, why is 2 squared and why is 5 a divisor?

terenzreignz (terenzreignz):

It was arbitrary.... I mean come on... \[\Large 2.4 = \frac{24}{10}= \frac{3\times 2\times2\times 2}{2\times 5}\] do the cancellations yourself.

OpenStudy (anonymous):

sorry,, it wasn't taught to us,

OpenStudy (anonymous):

okay i get it

terenzreignz (terenzreignz):

It's something you're already supposed to know... and it's called resourcefulness XD You needed to find a way to reduce 2.4 into an expression with numbers you can actually use.. again, 2's, 3's, 5's and 7's are the only logs given to you, so use those.

OpenStudy (anonymous):

log2+log4=log3+2log2-log5

terenzreignz (terenzreignz):

log 3 + 2log2 - log 5 correct log 2 + log 4....lacks the -log 5 :P

terenzreignz (terenzreignz):

So your final answer?

OpenStudy (anonymous):

why does it lack -log5?

terenzreignz (terenzreignz):

kako... kulang ng -log 5 :P log3 + log4 - log5 = log3 + 2log2 - log 5 :D

OpenStudy (anonymous):

bakit nga, ehh log(2.4) lang

terenzreignz (terenzreignz):

/hai/ \[\Large \log(2.4) = \log\left(\frac{3\cdot 4}{5}\right)=\log(3)+\log(4) - \log(5) \]okay? :P

OpenStudy (anonymous):

okaaaay,,

OpenStudy (anonymous):

akala ko kasi kasama si log(2.4)

OpenStudy (anonymous):

next?

terenzreignz (terenzreignz):

you already got it :/ log(3) + 2log(2) - log(5)

OpenStudy (anonymous):

that's the answer??

terenzreignz (terenzreignz):

Well, yes, actually :P Just replace the log(3), log(2) and log(5) with their actual values and simplify ^_^

OpenStudy (anonymous):

so i need calculator huh?

terenzreignz (terenzreignz):

not really... but go ahead :D

OpenStudy (anonymous):

how to simplify w/o calculator?

terenzreignz (terenzreignz):

simple. wag tamarin :P 0.4471 + 2(0.3010) - 0.699

OpenStudy (anonymous):

aiyt,, ganun lang,,ahahahah